# Difference of surface effects in the inclusive and reactions

### v1

@@ Abstract
Background: Previous studies have revealed surface effects in inclusive and reactions. These findings have contributed significantly to the improvement of nuclear data evaluation. However, surface effect in inclusive reaction has hardly been investigated.
Purpose: The purpose of this study is to investigate the difference of the surface effects in the and reactions. We also aim to clarify the origin of the difference.
@@
Results: In the region with small energy transfer, the radial distribution of the energy spectra for the reaction is found to be shifted toward the outer region of the nucleus more than that of the reaction. Furthermore, the calculations ignoring the nuclear absorption show that the radial distributions in the two reactions are nearly identical.
Conclusion: The surface effect in the reaction is more pronounced than in the reaction. A significant cause of the difference is that deuteron undergoes stronger absorption by the nucleus than proton.
@@ I INTRODUCTION
Phenomenological models have played an important role in the nuclear data evaluation for nucleon-induced reactions [1, 2, 3, 4, 5]. For the calculation of pre-equilibrium processes, which are prominent at incident energies above about 10 MeV, the exciton model [6] has been widely used. This phenomenological model has achieved great success in combination with global parameterization based on analyses of a large number of experimental data for reactions [7, 8]. It has been argued in these analyses that the inclusion of a surface localization effect is important in reproducing the shape of experimental energy spectrum [7, 9]. This effect is taken into account by introducing corrections to the density of states in the exciton model. Note that the importance of the surface effect is supported not only by the comparison with experimental data but also by the microscopic theoretical analysis [10] using the semiclassical distorted wave (SCDW) model, which has no free adjustable parameter [11, 12, 13, 14, 15].
@@
Evaluation activities of deuteron nuclear data have gradually begun [1, 16, 4], mainly motivated by the development of deuteron accelerator-based intensive neutron sources [17, 18]. In the evaluation of deuteron nuclear data, phenomenological models [19, 20] using state densities similar to those in the exciton model are often employed to calculate the components of inelastic scattering to continuum states. The evaluation of the inelastic scattering is important since it affects the transport of deuterons in a material. Therefore, as is the case with the exciton model, whether the surface effect should be taken into account in the state densities used in the above inelastic scattering model, and if so, to what extent, is an important factor in the deuteron nuclear data evaluation. However, since experimental data for reactions are very limited, there is concern that making this judgment based on a comparison with specific experimental data may give unphysical results under other conditions. A deep and comprehensive understanding for surface effects in reactions is required.
Under these circumstances, in Ref. [21], the one-step SCDW model [11] has been successfully applied to the description of inclusive reactions. The SCDW model has the property which describes the inclusive cross section as an incoherent integral of the contributions at individual collision points [22, 10]. Therefore, the model is expected to be useful for investigating the peripherality of the reaction. Findings from the microscopic reaction model analysis will provide guidance and support for the consideration of the surface effect in the phenomenological deuteron inelastic scattering models.
@@
The purpose of this study is to investigate the surface effect in the reaction and to clarify the difference from that in the reaction. We also aim to clarify the cause of the difference. The above analyses are performed utilizing the above-mentioned property of the SCDW model. Exploring the difference in surface effects between the two reactions are of interest from the view point not only of nuclear data evaluation but also of fundamental nuclear physics. Furthermore, we apply the findings from the SCDW analysis to the implementation of the surface effect into the phenomenological deuteron inelastic scattering model. This is intended to show how much microscopic reaction theory contributes to improving phenomenological models.
@@
The construction of this paper is as follows. In Sec. II we briefly describe the one-step SCDW model for the inclusive and reactions. We also explain here the phenomenological inelastic scattering model. In Sec. III we compare the calculated energy spectrum of the inclusive and reactions with experimental data and clarify the difference of the surface effects between the two reactions. Applications for improving the phenomenological model also be conducted here. Finally, a summary is given in Sec. IV.
@@ II.1 SCDW model
We briefly describe the inclusive and reactions with the one-step SCDW model. The double differential cross section for the energy and the solid angle of the emitted particle is expressed with
@@
where the subscript denotes the label of proton and neutron in the target. is the momentum of the nucleon in the target in the initial state. The local Fermi momenta of nucleon and are denoted by and , respectively. is the free scattering cross section determined by the local scattering angle and the local scattering energy between and the nucleon in the target. is the incident energy of and is the kinetic energy of the nucleon in the target nucleus in the initial (final) state.
@@
By applying the local Fermi gas model (LFG) to the initial and final nuclear single-particle states, is given by [13]
where and are atomic and neutron numbers of the target nucleus, respectively. LFG can takes into account the spread of nucleon density on the nuclear surface because the local Fermi momentum depends on . To clarify the surface effect, we also consider the Fermi gas model (FG), which give a uniform Fermi momentum distribution. When FG is used instead of LFG, and become step functions with respect to in Eq. (2). On the other hand, the Fermi momentum in LFG decreases smoothly as increases, as shown later in Fig. 1.
@@ II.2 Phenomenological model by Kalbach
As a phenomenological model for continuum deuteron inelastic scattering, we adopt the model proposed by Kalbach [28]. Although this model was originally developed for -particle-induced reactions, our previous study has shown that the model is applicable to deuteron-induced ones [19].
According to Ref. [28], the energy spectrum of the deuteron inelastic scattering with a nucleon particle-hole pair creation is calculated by:
where is a normalization constant determined from fitting experimental energy spectrum, is the deuteron total reaction cross section at scattering energy , and and are the incident and emitted deuteron energies, respectively. is obtained from the optical model calculation with the global potential by An-Cai [24]. is the spin of the emitted deuteron, respectively. denotes the mass number of the target, and represents the excitation energy of residual nucleus. The final state density is expressed as follows:
where , , and and are the numbers of neutrons and protons in the target. The setting of values for and follows the original paper by Kalbach [28]. The squaring of and corresponds to considering the - states of nucleons. The finite well function with the well depth is defined for the - states as follows:
where is the unit step function.
A simple method to include surface effects in the exciton model was proposed in Ref. [7], and has been successfully applied to the analysis of reactions [9, 8]. In the method, the surface effects are taken into account by setting the small value of for the =1 state. In this work we apply this method to Eq. (6). In other words, we set in Eq. (8) smaller than the typical value of 38 MeV associated to the Fermi energy. The details of setting the value of are described later in Sec. III.3.
@@ III RESULTS AND DISCUSSION
III.1 Effect of momentum distribution of nucleons on energy spectrum
@@
In FG, no reaction is allowed in fm because is zero in the region. On the other hand, in LFG, reactions are allowed beyond the boundary because LFG incorporates the effect of diffuseness of the nucleus. Therefore, the difference of the SCDW calculations with LFG and FG can be tied to surface effects.
Figure 2 shows the energy spectra of the and reactions for several incident energies and targets. The experimental data are taken from Refs. [29, 30]. In order to make the velocity of projectile same, the incident energies per nucleon in Fig. 2(e) and (f) are the same as those in (b) and (c), respectively. Note that these results are calculated with SCDW which has no free adjustable parameter.
@@
By comparing the energy spectra calculated with LFG and FG, we can see that consideration of the surface effects is necessary to reproduce experimental data in both of the and reactions. The surface effects do not depend so much on the incident energy and target nucleus in both reactions. On the other hand, the reactions have stronger dependence than that of the reactions.
III.2 Difference of radial distributions of and reactions
To analyze the difference of the surface effects between the and reactions in more detail, we discuss the radial distribution of the energy spectra.
@@ III.2 Difference of radial distributions of and reactions
As illustrated in Fig. 3, for a given , the peak position is almost the same for the and reactions and the peak positions shift outwards as decreases. In contrast, the dependence of the peak heights of the two reactions is opposite. The peak height of the reaction becomes smaller as decreases, while that of the reaction becomes larger as decreases.
To clearly show how this significant difference affects the surface effects of these reactions, we present in Fig. 4 the integrated values of in the range of 5 15 MeV. The solid and dashed lines represent the calculated values of the and reactions, respectively. Each of them is normalized so that the integrated value over is unity. It is seen that the radial peak position is about 4.3 fm for the reaction. On the other hand, for the reaction, the peak position is about 6.0 fm and the large proportion of reaction occurs in the outer region of the nucleus. This clearly shows that the surface effect is more significant for the reaction.
Next, we discuss the cause of the difference in the surface effects between the two reactions. Figure 5 is the same as Fig. 4 but for the calculations ignoring the absorption of the protons and deuterons by the nucleus; the imaginary part of the optical potential is set to zero in the calculations. When we ignore the absorption, the radial distributions of the two reactions are almost identical. This means that the stronger surface effect in the reaction than in the reaction is attributed to the strong absorption of deuteron by the nucleus. We can also see that, in Fig 3, the strong absorption of the deuterons suppresses the reaction in the inner region of the nucleus in the than in the reaction.
@@ III.3 Application to phenomenological model
Next, we apply the findings obtained from the previous sections to the implementation of the surface effect into the phenomenological model. Figure 6 illustrates the comparison of the experimental data and the energy spectra calculated with the Kalbach model presented in Sec II.2. The experimental data are obtained from Ref. [30].
@@
Three lines are shown in the figure. The solid line is the result of the calculation in which the well depth is optimized to reproduce the experimental spectrum. The optimised value of is 10 MeV. The normalization factor is adjusted so that the peak of the spectrum matches the experimental data. This adjustment holds true also for the two calculations below. The dashed line is the result with the parameterization optimized for proton-induced reactions [8]. In Ref. [8], the analyses using the exciton model were performed for reactions up to 200 MeV for various targets, and the different parameterizations for neutron- and proton-induced reactions were given based on the results. The value of in the case of proton-induced reactions is given as follows:
@@
where and are the incident energy and the target mass, respectively. In the case shown in Fig. 6, Eq. (10) gives about 25 MeV as the value of . The dotted line is the calculation with a typical well depth of 38 MeV. Using this value of corresponds to not considering the surface effects.
@@
As shown in the figure, the calculation not considering the surface effects does not reproduce the shape of the experimental spectrum. Note that the components below 30 MeV are mainly due to the pre-equilibrium process for other states than - and the compound nucleus process. The large component slightly below 80 MeV is attributed to inelastic scattering to low-lying discrete levels and elastic scattering. These components are not considered in the Kalbach model and are thus outside the scope of the present discussion.
@@
In terms of the calculations considering the surface effects, the use of the optimized value for proton-induced reactions improves the agreement with experimental data, but underestimation in the high-energy region is still seen. On the other hand, when the value of is decreased as our SCDW analysis suggests and the value is set to 10 MeV, the calculation results reproduce the experimental energy spectrum well over almost the entire energy range to be considered in the Kalbach model. These results indicate that parameter optimization of the phenomenological Kalbach model has been justified also by theoretical analysis with the microscopic SCDW model.

### v2

# Difference in peripherality of the inclusive and reactions and its implications for phenomenological reaction model

@@ Abstract
Background: Previous studies have revealed the importance of introducing surface correction into a phenomenological model for inclusive and reactions. These findings have contributed significantly to the improvement of nuclear data evaluation. However, the necessity for the surface correction in an inclusive reaction has hardly been investigated.
Purpose: The purpose of this study is to investigate the difference in the peripherality of the and reactions by a theoretical analysis using a quantum mechanical model, and to obtain a theoretical basis on the surface correction in the reaction.
@@
Results: The radial distribution of the energy spectra for the reaction is shifted toward the outer region of the nucleus compared to the reaction. Based on this finding, we consider a larger surface correction into a phenomenological model for the reaction than that for the reaction, and calculated values reproduce the experimental spectra well.
Conclusion: The peripherality of the reaction is more prominent than that of the reaction. The stronger surface correction thus should be introduced for the reaction than for the reaction.
@@ I INTRODUCTION
Phenomenological models have played an important role in the nuclear data evaluation for nucleon-induced reactions [1, 2, 3, 4, 5]. For the calculation of pre-equilibrium processes, which are prominent at incident energies above about 10 MeV, the two-component exciton model [6] has been widely used. This phenomenological model has achieved great success in combination with global parameterization based on analyses of a large number of experimental data for reactions [7, 8, 9]. It has been argued in these analyses that the correction related to surface localization is important in reproducing the shape of experimental energy spectrum [7, 8]. Note that we will refer to the above correction as “surface correction” in this paper even though the correction has been referred to as “surface effects” in previous literature including Refs. [7, 8, 9]. This is because the “surface effects” is actually interpreted as a correction within the phenomenological model, as explained below.
The surface correction is a simple means to effectively incorporate the peripherality of the reaction into the exciton model by restricting a hole degree of freedom with respect to the excitation energy. The original exciton model [10] assumes that the hole states could extend to infinite depth. This model was later modified by considering a finite well depth in the density of states [11]. The modification restricts the region where the hole states are generated to a more realistic range. In the surface correction, the finite well depth is made shallower only in the one-hole state, i.e., the state where the number of collision is small as discussed in Ref. [7]. This is an idea based on the following two matters in the exciton model that does not have a radial dependence: (i) reactions with fewer collisions would be more likely to occur around the nuclear surface, (ii) nuclear potential is shallower around the nuclear surface than the interior. It is worth mentioning that in Refs. [12, 13], it is indicated that the first collision in reactions is localized on the nuclear surface using a semiclassical approach. Furthermore, the peripherality of the reaction is clarified by the theoretical analysis [14] using the semiclassical distorted wave (SCDW) model, which has no free adjustable parameter [15, 16, 17, 18, 19].
@@
On another front, evaluation activities of deuteron nuclear data have gradually begun [1, 20, 4], mainly motivated by the development of deuteron accelerator-based intensive neutron sources [21, 22]. In the evaluation of deuteron nuclear data, phenomenological models [23, 24] using state densities similar to those in the exciton model are often employed to calculate the components of direct inelastic scattering to continuum states. These components correspond to inelastic scattering processes involving excitation of a nucleon particle-hole pair in the target nucleus. The evaluation of the inelastic scattering is important since it affects the transport of deuterons in a material. Therefore, as is the case with the exciton model, whether the surface correction should be taken into account in the state densities used in the above inelastic scattering model, and if so, to what extent, is an important factor in the deuteron nuclear data evaluation. However, since experimental data for the reactions are very limited, there is concern that making this judgment based on a comparison with specific experimental data may give unphysical results under other conditions. A theoretical basis for the surface correction of the reaction founded on a deep and comprehensive understanding of the peripherality of the reaction is required.
Under these circumstances, in Ref. [25], the one-step SCDW model [15] has been successfully applied to describe the inclusive reactions. This SCDW model has two noteworthy features. First, the model possesses the property of representing the inclusive cross section as an incoherent integral of contributions at individual collision points [26, 14]. This property allows us to analyze the peripherality of the inclusive reaction, i.e., to investigate where and to what extent the reaction occurs in the nucleus. The second is the adoption of local Fermi gas (LFG) to the single-particle states. Although LFG would be unrealistic for describing a specific nuclear state, it can reasonably describe the overall response of a nucleus, involving many initial and final states. Note that, in Ref [18], the SCDW model adopts the Wigner transform of one-body density matrices calculated with a single-particle state model for nuclei instead of LFG. On the other hand, we use LFG for reducing numerical tasks in this work.
@@
The purpose of this study is to investigate the peripherality in the inclusive reaction and the cause of the difference from that in the inclusive reaction. The analyses are performed utilizing the above-mentioned property of the SCDW model. Exploring the difference in the peripherality of the two reactions is of interest from the viewpoint not only of nuclear data evaluation but also of fundamental nuclear physics. Furthermore, we apply the findings from the SCDW analysis to the implementation of the surface correction into the phenomenological deuteron inelastic scattering model.
@@
The construction of this paper is as follows. In Sec. II we briefly describe the one-step SCDW model for the inclusive and reactions. We also explain the phenomenological inelastic scattering model. In Sec. III we compare the calculated energy spectrum of the inclusive and reactions with experimental data and clarify the difference of the peripherality between the two reactions. Improvements of the phenomenological model are also suggested. Finally, a summary is given in Sec. IV.
@@ II.1 SCDW model
We briefly describe the inclusive and reactions with the one-step SCDW model. The double differential cross section for the energy and the solid angle of the emitted particle is expressed with [25]
@@
where is the momentum of the nucleon of the target in the initial state. represents the local Fermi momenta of nucleon. is the free scattering cross section determined by the local scattering angle and the local scattering energy between and the nucleon of the target. The incident energy of is represented by and is the kinetic energy of the nucleon of the target nucleus in the initial (final) state. Equations (1) and (2) are the same as Eqs. (19) and (20) of Ref. [25], respectively, and the derivation of the equations is discussed in detail in Ref. [25].
From Eq. (1), one can see that the inclusive cross section is described by an incoherent integral of . This description is made possible by the short-ranged property of the kernel, Eq. (2.8) of Ref. [15], which is realized when a large number of s.p. states are involved; see Fig. 11 of Ref. [15]. In other words, the interference between the transition through different interacting points disappears in the situation. Thus, the projectile-nucleus cross section is given by integrating a product of the probability of the projectile reaching the point , that of the ejectile being emitted from , and the averaged cross section of the elementary process at . For more details, readers are referred to Ref. [15].
By applying the local Fermi gas model (LFG) to the initial and final nuclear single-particle states, is given by [15]. Therefore, LFG can takes into account the spread of nucleon density on the nuclear surface. To clearly show the contribution of the nuclear surface to the inclusive cross section, we also consider the Fermi gas model (FG), which gives a uniform Fermi momentum distribution. When FG is used instead of LFG, becomes step functions with respect to in Eq. (2). On the other hand, the Fermi momentum in LFG decreases smoothly as increases, as shown later in Fig. 1.
@@ II.2 Phenomenological model by Kalbach
As a phenomenological model for continuum deuteron inelastic scattering, we adopt the model proposed by Kalbach [32]. Although this model was originally developed for -particle-induced reactions, the model itself is considered to be applicable to deuteron-induced ones [23]. Note that what is calculated in this model is the direct process component, which is not included in the exciton model.
According to Ref. [32], the energy spectrum of the direct deuteron inelastic scattering with a nucleon particle-hole pair creation is calculated by:
where is a normalization constant determined from fitting experimental energy spectrum, is the deuteron total reaction cross section at scattering energy , and and are the incident and emitted deuteron energies, respectively. According to Ref. [32], the overall normalization constant can vary in the range of a factor of 3 depending on the target. As shown later in Sec. III.3, in this study, is adjusted so that the peak of the spectrum matches available experimental data. is obtained from the optical model calculation with the global potential by An-Cai [28]. is the spin of the emitted deuteron. denotes the mass number of the target, and represents the effective excitation energy of residual nucleus considering the paring effect [33]. The final state density is expressed as follows:
where , , and and are the numbers of neutrons and protons in the target. The setting of values for and follows the original paper by Kalbach [32]. The squaring of and corresponds to considering the - states of nucleons. The finite well function with the well depth is defined for the - states as follows [11]:
where is the unit step function. Note that the finite well function is not taken into account in the original Kalbach model [32] and the first application of the model to the deuteron-induced reaction [23]. The introduction of the finite well function enables us to consider the surface correction as shown below.
As mentioned in Sec. I, a simple method to include the peripherality of the nuclear reaction in the exciton model was proposed in Ref. [7], and has been successfully applied to the analysis of reactions [8, 9]. In the method, the peripherality is effectively taken into account by fixing the small value of for the =1 state only. In this work we apply this method to Eq. (4). In other words, we set in Eq. (6) smaller than the typical value of 38 MeV associated to the Fermi energy. By introducing is replaced by smaller density of states when . Therefore, giving a smaller V corresponds to making the reaction more likely occur in the peripheral region of the nucleus, based on the fact that the nuclear potential is shallower around the nuclear surface than in the interior. The details of fixing the value of are described later in Sec. III.3.
Following the idea in Kalbach’s original paper [32], the values calculated with the inelastic scattering model are treated as the direct component in this study. Therefore, we do not consider in the model the reduction of the composite nucleus production cross section caused by other direct processes, such as the breakup reactions. The use of the total reaction cross section obtained from the optical model as in Eq. (4) reflects the above idea. The interference with other direct processes is not considered.
@@ III RESULTS AND DISCUSSION
III.1 Influence of nuclear surface on energy spectrum
@@
In FG, no reaction is allowed in fm because is zero in the region. On the other hand, in LFG, reactions are allowed beyond the boundary because LFG incorporates the diffuseness around the nuclear surface. Therefore, the difference between the SCDW calculation with LFG and that with FG can be tied to the influence of considering the nuclear surface.
Figure 2 shows the energy spectra of the and reactions for several incident energies and targets. The horizontal axis is the energy transfer . The solid (dashed) lines are calculated results with LFG (FG). The experimental data are taken from Refs. [34, 35]. In order to make the velocity of projectile same, the incident energies per nucleon in Fig. 2(e) and (f) are the same as those in (b) and (c), respectively. Note that these results are calculated with SCDW which has no free adjustable parameter.
@@
By comparing the energy spectra calculated with LFG and FG, we can see that consideration of the nuclear surface is necessary to reproduce experimental data in both the and reactions. The importance of considering the nuclear surface does not depend so much on the incident energy and target nucleus in both reactions. We can also find that the influence of the nuclear surface gets larger as decreases. This result implies that the peripherality of the reactions becomes stronger as decreases. Moreover, the dependence is more prominent in the reactions than the reactions.
III.2 Difference in peripherality of and reactions
To analyze the difference in the influence of nuclear surface between the and reactions in more detail, we discuss the radial distribution of the energy spectra.
@@ III.2 Difference in peripherality of and reactions
As shown in Fig. 3, for a given , the peak position is almost the same for the and reactions and the peak positions shift outwards as decreases. In contrast, the dependence of the peak heights of the two reactions is opposite. The peak height of the reaction becomes smaller as decreases, while that of the reaction becomes larger as decreases.
To clearly show how this significant difference affects the peripherality of these reactions, we present in Fig. 4 the integrated values of in the range of 5 15 MeV. The solid and dashed lines represent the calculated values of the and reactions, respectively. Each of them is normalized so that the integrated value over is unity. It is seen that the radial peak position is about 4.3 fm for the reaction. On the other hand, for the reaction, the peak position is about 6.0 fm and a large proportion of the reaction occurs in the outer region of the nucleus. This result clearly shows that the reaction has a stronger peripherality than the reaction.
Next, we discuss the cause of the difference in the peripherality of the two reactions. Figure 5 is the same as Fig. 4 but for the calculations ignoring the absorption of the protons and deuterons by the nucleus; the imaginary part of the optical potential is set to zero in the calculations. When we ignore the absorption, the radial distributions of the two reactions are almost identical. Therefore, we can see that the stronger peripherality of the reaction than that of the reaction is attributed to the strong absorption of deuteron by the nucleus. We can also confirm that in Fig 3, the strong absorption of the deuteron suppresses the reaction in the inner region of the nucleus in the than in the reaction.
@@ III.3 Application to phenomenological model
Next, we apply the findings obtained from the previous sections to the implementation of the surface correction into the phenomenological model. Figure 6 shows the comparison of the experimental data and the energy spectra calculated with the Kalbach model presented in Sec. II.2. The horizontal axis represents the energy transfer . The experimental data are obtained from Ref. [35].
@@
Four lines are shown in the figure. The solid line is the result of the calculation in which the well depth is optimized to reproduce the experimental spectrum. The optimised value of is 10 MeV. The normalization factor is adjusted so that the peak of the spectrum matches the experimental data. This adjustment holds true also for the three calculations below. The dashed line shows the result of using 17 MeV as the value of . In Ref. [8], a value of 17 MeV was proposed as the optimum value of for the reaction, independent of the incident energy and target nucleus. A local optimum value of MeV has also been derived in Ref. [8]. On the other hand, it is also discussed in Ref. [8] that the value of can vary depending on the background of the experiments being compared. We thus adopted the global optimum value of MeV, obtained from the analysis of several tens of experiments, as one of the targets for comparison. The dotted line is the result with the parameterization optimized for proton-induced reactions derived in Ref. [9]. In Ref. [9], the analyses using the exciton model were performed for reactions up to 200 MeV for various targets, and the different parameterizations for neutron- and proton-induced reactions were given based on the results. The value of in the case of proton-induced reactions is given as follows:
@@
where and are the incident energy and the target mass, respectively. In the case shown in Fig. 6, Eq. (8) gives about 25 MeV as the value of . On the other hand, Finally, the dash-dotted line is the calculation with a typical well depth of 38 MeV. Using this value of corresponds to not introducing the surface correction.
Note that in Ref. [23] we have performed an analysis of the double differential cross sections for the same target and incident energy at 20°. At such a forward angle, the components corresponding to giant resonance states were dominant in the region of MeV. On the other hand, the use of angle-integrated DX as in Fig. 6 makes it easier to discuss the direct inelastic scattering component calculated with the Kalbach model, since the giant resonance components make up a smaller fraction of the total in the above region.
@@
As shown in the figure, the calculation not introducing the surface correction does not reproduce the shape of the experimental spectrum. Note that the components above 50 MeV are mainly due to the pre-equilibrium process and the compound nucleus process. The large component below a few MeV is attributed to inelastic scattering to low-lying discrete levels and elastic scattering. These components are not considered in the Kalbach model and are thus outside the scope of the present discussion.
@@
In terms of the calculations incorporating the surface correction, the use of the optimized values for proton-induced reactions improves the agreement with experimental data, but underestimation in the high-energy region is still seen. On the other hand, when the value of is more decreased, the calculation results reproduce the experimental energy spectrum better over almost the entire energy range to be considered in the Kalbach model. As discussed in Sec. II.2, the smaller value of makes the reaction more likely occur in the peripheral region of the nucleus. Therefore, this result is consistent with the results of the present SCDW analysis. In other words, parameter optimization of the phenomenological Kalbach model has been justified also by theoretical analysis with the SCDW model.
In Ref. [7], it is argued that the magnitude of the surface correction can also vary depending on what excitation energy dependence is assumed for the single-particle level density. Meanwhile, an equispacing model that does not assume an excitation energy dependence of the single-particle level density is adopted in Refs. [9, 8], and the present study also uses the same equispacing model as presented in Eq. (5). It is worth mentioning that these comparisons thus are valid as a discussion on the surface correction.